The greatest angle of a cyclic quadrilatral is double the Least and difference of other two angles is 30 degree.find the angles

Grade:11

1 Answers

Ritesh Khatri
6 years ago
Let ABCD be quadrilateral.
Now Sum of opposite angles of a cyclic quadrilateral is equal to 180 degree.
therefore if angle A is greatest then C must be Least.
 
A + C = 180
Also A = 2C
 
Therefore angle C = 60 degrees and angle A = 120 degrees
 
Also B+D = 180 
Let B – D = 30
 
Hence angle B = 105 degrees and Angle D = 75degree
Therefore Angles are 30 , 60 , 105 and 75 degrees